1) Diana has a total of 100 dimes and quarters.If the total value of the coins is $13.75, how many quarters does she have?
Answer by mananth(16946)

quarters value= 25x
dimes Value=10 (100-x)
Value= 1375
25 x -10 x + 1000 = 1375
15 x + 1000 = 1375
15 x = 1375 -1000
15 x = 375
/ 15
x= 25 quarters 625
75 dimes 750